Water moves through the entire diameter of 3 cm at a speed of 3 m / s. In one part of the water it enters a narrowed part whose speed is 15m / s. Determine the cross-sectional diameter of the narrowed part of the pipe. (r2 = 0.67cm)


1
Expert's answer
2021-04-05T11:10:40-0400

"A_1v_1=A_2v_2\\0.25\pi d_1^2v_1=0.25\pi d_2^2v_2\\
(3)^2(3)=d_2^2(15)\\d_2=1.34\ cm"